OpenStudy (anonymous):

Train A and B are traveling in the same direction on parallel tracks. Train A is traveling at 100 miles per hour and train B is traveling at 120 miles per hour. Train A passes a station at 10.10 A.M. If train B oases the same station at 10:25 A.M., at what time will train B catch up to train A?

OpenStudy (anonymous):

time is only an illusion of relative motion

OpenStudy (lgbasallote):

does it say anything about the two trains' starting at the same position, or anything like that?

OpenStudy (anonymous):

no thats all the info

OpenStudy (dumbcow):

At the time train B catches up with train A, they would have traveled same distance from station except it would have taken Train B 15 min less time to get there. **Note: 15 min = 0.25 hr d = r*t 120(t-.25) = 100t 120t - 30 = 100t 20t = 30 t = 1.5 It took 1.5 hrs from time Train A passed station for train B to catch up, or at 11:40